# Mirror-Neuron Patterns in AI Alignment

### v1

@@ 1.3 Scope and Objectives
    Propose a theoretical framework to explain the emergence of mirror neuron patterns, incorporating concepts like neuronal economy, mutual dependency, and the Veil of Ignorance.
@@ 2.5 Empathy in Artificial Intelligence
2.5.1 Shared Representations and Neuronal Economy
When models are excessively large or unconstrained, they can allocate resources to memorize individual states, bypassing the need for shared representations [26, 27]. Conversely, appropriately scaled ANNs, learn reusable patterns that span multiple scenarios, requiring the network to economize resources and avoid overfitting to specific conditions [28, 29]. This dynamic, which we term neuronal economy, is a vital precursor to the empathic-like behaviors explored in this research.
@@ 2.7 Summary
The importance of this is underscored by current debates on risk, which highlight the limitations of external constraints and the urgent need for AI systems capable of intrinsic alignment with human values to mitigate catastrophic risks effectively. By examining how mirror neuron patterns might emerge in simple ANNs and integrate with concepts like neuronal economy, agent dependency, and the Veil of Ignorance, this dissertation lays the groundwork for embedding intrinsic ethical motivations in AI. In doing so, it moves beyond external rule sets and top-down constraints, and explores whether affective empathy-like processes can serve as an internal moral compass, guiding AI toward safer and more ethical behavior.
@@ 3.2 Emergent Factors from Early Experiments
While agent dependency ($D$) was initially believed to be the primary factor, further experiments revealed that mirror neuron pattern emergence was not solely dependent on $D$. Two additional factors – Neuronal Economy and Veil of Ignorance – emerged as significant contributors. These factors are explored in detail below.
3.3 Neuronal Economy
Neuronal Economy describes the efficiency with which an ANN utilizes its resources – Signal Complexity ($S$), Model Capacity ($M$), and Error ($E$) – to generalize and form shared neural representations. It is captured by:
@@ 3.3 Neuronal Economy
This defines the Neuronal Economy Function, which balances:
@@
If $M$ (Model Capacity) is too low relative to $S$ (Signal Complexity), the network struggles to capture the intricacies of the signal and fails to produce meaningful generalizations, resulting in high error. However, Neuronal Economy is not merely a bias-variance tradeoff but reflects a distinct phenomenon observed in our experiments. When $S$ is too low relative to $M$, the network overfits, effectively creating a lookup table instead of generalizing. This condition corresponds to poor Neuronal Economy, as the network fails to develop shared neural representations despite achieving low error.
An optimal Neuronal Economy exists within a range where S/M and E are balanced, enabling the network to generalize across scenarios and form shared neural representa- tions.
@@ 3.5 Proportionality of Factors Influencing Mirror Neuron Emergence
The probability $P$ of mirror neuron emergence is hypothesized to be proportional to the combined influence of two key factors: Neuronal Economy and Self/Other Relation. Formally:
@@
  • $f\left(\dfrac{S}{M},\,E\right)$, the Neuronal Economy Function, captures the balance between Signal Complexity ($S$), Model Capacity ($M$), and Error ($E$). This function ensures the network maintains balance, avoiding excessive complexity or underutilized capacity, supporting the emergence of shared neural representations. While these are critical for generalization, they do not inherently involve agency or relational dynamics.
@@ 4.1 Experimental Goals
While many neural networks can form shared representations, these experiments are designed not only to encourage generalization (as described by the Neuronal Economy, $f\left(\tfrac{S}{M},\,E\right)$) but also to introduce conditions of agent dependency and identity uncertainty (captured by the Self/Other Relation Function, $d\left(D,\,I\right)$). This ensures that we test for the emergence of patterns that go beyond shared representations and approach the relational dynamics g(D,I)g(D,\,I), hypothesized to underlie mirror neuron-like behavior.
@@ 4.3.2 Training Process
By adjusting the complexity of the Frog and Toad environment and the ANN’s capacity, we effectively tested different $\tfrac{S}{M}$ conditions. Combined with early stopping and error minimization, these variations allowed us to infer how changes in these parameters affected the network’s Neuronal Economy $f\left(\tfrac{S}{M},E\right)$ and, by extension, its tendency to form shared neural representations.
@@ 7.1.1 Awareness of Self and Other in Neural Networks
While these "self-other" distinctions are task-bound abstractions rather than genuine biological awareness, the observed patterns align with our theoretical constructs. The network’s ability to form shared representations corresponds to achieving a balanced Neuronal Economy ($f(\tfrac{S}{M},E)$), while conditions fostering self/other relations reflect the influence of Agent Dependency and the Veil of Ignorance ($g(D,I)$). Thus, the emergent behaviors directly support the proposed theoretical framework.
@@ 7.2.3 Theoretical Framework
  • $f\left(\tfrac{S}{M},\,E\right)$ represents the Neuronal Economy Function, capturing how the balance between signal complexity ($S$) and model capacity ($M$), along with error ($E$), fosters generalization and shared neural representations.
@@ 7.3 Conclusion
This dissertation demonstrates that mirror neuron patterns can emerge in simple artificial neural networks (ANNs) and suggests how these patterns might contribute to ethical AI alignment. Through a novel experimental framework combining neuronal economy $f\!\bigl(\tfrac{S}{M},E\bigr)$ and a self/other relation $g(D,I)$, we show that mirror neuron-like representations arise when networks are appropriately scaled to input signals, and where self-experienced and observed conditions involve agent dependency alongside a limit on self/other differentiation. Analysis of inter-layer connections reveals how certain neurons integrate empathic signals from mirror neuron candidates with agent-differentiating cues. These findings provide compelling evidence that the network has modeled another’s distress as if it were its own, illustrating how internal simulation can support prosocial action.
@@
This work opens new possibilities at the intersection of artificial intelligence and cognitive science, offering a promising direction for advances in AI ethics, neuroscience, and the development of cooperative artificial systems. Future research should further refine the neuronal economy function $f\!\bigl(\tfrac{S}{M},E\bigr)$ and the self/other relation function $g(D,I)$, quantifying their precise forms and testing their applicability in more complex, dynamic environments and larger-scale models. Such intrinsic alignment mechanisms, like those demonstrated here, might scale to address existential risks and promote long-term safety in advanced AI systems.

### v2

@@ 1.3 Scope and Objectives
    Propose a theoretical framework to explain the emergence of mirror neuron patterns, incorporating concepts like neural economy, mutual dependency, and the Veil of Ignorance.
@@ 2.5 Empathy in Artificial Intelligence
2.5.1 Shared Representations and Neural Economy
When models are excessively large or unconstrained, they can allocate resources to memorize individual states, bypassing the need for shared representations [26, 27]. Conversely, appropriately scaled ANNs, learn reusable patterns that span multiple scenarios, requiring the network to economize resources and avoid overfitting to specific conditions [28, 29]. This dynamic, which we term Neural Economy, is a vital precursor to the empathic-like behaviors explored in this research.
@@ 2.7 Summary
The importance of this is underscored by current debates on risk, which highlight the limitations of external constraints and the urgent need for AI systems capable of intrinsic alignment with human values to mitigate catastrophic risks effectively. By examining how mirror neuron patterns might emerge in simple ANNs and integrate with concepts like neural economy, agent dependency, and the Veil of Ignorance, this dissertation lays the groundwork for embedding intrinsic ethical motivations in AI. In doing so, it moves beyond external rule sets and top-down constraints, and explores whether affective empathy-like processes can serve as an internal moral compass, guiding AI toward safer and more ethical behavior.
@@ 3.2 Emergent Factors from Early Experiments
While agent dependency ($D$) was initially believed to be the primary factor, further experiments revealed that mirror neuron pattern emergence was not solely dependent on $D$. Two additional factors – Neural Economy and Veil of Ignorance – emerged as significant contributors. These factors are explored in detail below.
3.3 Neural Economy
Neural Economy describes the efficiency with which an ANN utilizes its resources – Signal Complexity ($S$), Model Capacity ($M$), and Error ($E$) – to generalize and form shared neural representations. It is captured by:
@@ 3.3 Neural Economy
This defines the Neural Economy Function, which balances:
@@
If $M$ (Model Capacity) is too low relative to $S$ (Signal Complexity), the network struggles to capture the intricacies of the signal and fails to produce meaningful generalizations, resulting in high error. However, neural economy in this context is not merely a bias-variance tradeoff but reflects a distinct phenomenon observed in our experiments. When $S$ is too low relative to $M$, the network overfits, effectively creating a lookup table instead of generalizing. This condition corresponds to poor neural economy, as the network fails to develop shared neural representations despite achieving low error.
An optimal neural economy exists within a range where S/M and E are balanced, enabling the network to generalize across scenarios and form shared neural representations.
@@ 3.5 Proportionality of Factors Influencing Mirror Neuron Emergence
The probability $P$ of mirror neuron emergence is hypothesized to be proportional to the combined influence of two key factors: Neural Economy and Self/Other Relation. Formally:
@@
  • $f\left(\dfrac{S}{M},\,E\right)$, the Neural Economy Function, captures the balance between Signal Complexity ($S$), Model Capacity ($M$), and Error ($E$). This function ensures the network maintains balance, avoiding excessive complexity or underutilized capacity, supporting the emergence of shared neural representations. While these are critical for generalization, they do not inherently involve agency or relational dynamics.
@@ 4.1 Experimental Goals
While many neural networks can form shared representations, these experiments are designed not only to encourage generalization (as described by the Neural Economy function, $f\left(\tfrac{S}{M},\,E\right)$) but also to introduce conditions of agent dependency and identity uncertainty (captured by the Self/Other Relation Function, $d\left(D,\,I\right)$). This ensures that we test for the emergence of patterns that go beyond shared representations and approach the relational dynamics g(D,I)g(D,\,I), hypothesized to underlie mirror neuron-like behavior.
@@ 4.3.2 Training Process
By adjusting the complexity of the Frog and Toad environment and the ANN’s capacity, we effectively tested different $\tfrac{S}{M}$ conditions. Combined with early stopping and error minimization, these variations allowed us to infer how changes in these parameters affected the network’s neural economy $f\left(\tfrac{S}{M},E\right)$ and, by extension, its tendency to form shared neural representations.
@@ 7.1.1 Awareness of Self and Other in Neural Networks
While these "self-other" distinctions are task-bound abstractions rather than genuine biological awareness, the observed patterns align with our theoretical constructs. The network’s ability to form shared representations corresponds to achieving a balanced Neural Economy ($f(\tfrac{S}{M},E)$), while conditions fostering self/other relations reflect the influence of Agent Dependency and the Veil of Ignorance ($g(D,I)$). Thus, the emergent behaviors directly support the proposed theoretical framework.
@@ 7.2.3 Theoretical Framework
  • $f\left(\tfrac{S}{M},\,E\right)$ represents the Neural Economy Function, capturing how the balance between signal complexity ($S$) and model capacity ($M$), along with error ($E$), fosters generalization and shared neural representations.
@@ 7.3 Conclusion
This dissertation demonstrates that mirror neuron patterns can emerge in simple artificial neural networks (ANNs) and suggests how these patterns might contribute to ethical AI alignment. Through a novel experimental framework combining neural economy $f\!\bigl(\tfrac{S}{M},E\bigr)$ and a self/other relation $g(D,I)$, we show that mirror neuron-like representations arise when networks are appropriately scaled to input signals, and where self-experienced and observed conditions involve agent dependency alongside a limit on self/other differentiation. Analysis of inter-layer connections reveals how certain neurons integrate empathic signals from mirror neuron candidates with agent-differentiating cues. These findings provide compelling evidence that the network has modeled another’s distress as if it were its own, illustrating how internal simulation can support prosocial action.
@@
This work opens new possibilities at the intersection of artificial intelligence and cognitive science, offering a promising direction for advances in AI ethics, neuroscience, and the development of cooperative artificial systems. Future research should further refine the neural economy function $f\!\bigl(\tfrac{S}{M},E\bigr)$ and the self/other relation function $g(D,I)$, quantifying their precise forms and testing their applicability in more complex, dynamic environments and larger-scale models. Such intrinsic alignment mechanisms, like those demonstrated here, might scale to address existential risks and promote long-term safety in advanced AI systems.